For the system: I2(g) + Cl2(g) ↔ 2ICl(g) at equilibrium 0.310 moles of I2, 0.310 moles of Cl2 and 2.81 moles of ICl was found in a 1L vessel. 1 mole of ICl was then added to the vessel. Calculate the final equilibrium concentration of Cl2
Answer #1

Molarity of I2 = moles / volume ( in litres )

= 0.310 / 1

= 0.310 M

Molarity of Cl2 = 0.310 / 1 = 0.310

Molarity of ICl = 2.81/ 1 = 2.81

Kc = (ICl)^2 / ((I2) * (Cl))

= ( 2.81 )^2 / ( 0.310*0.310)

= 82.17

After adding 1 moles

ICl = 3.81

Now

I2 + Cl2 = 2 ICl

0.310 0.310 3.81 ( initially )

0.310+x 0.310+x 3.81-2x (equilibrium)

Kc = (ICl)^2 / ( I2*Cl2)

82.17 = (3.81-2x)^2 / ((0.310+x)*(0.310+x))

x = 0.0904

Cl2 = 0.310 + x = 0.310 + 0.0904 = 0.4004 M

Any query please comment

If satisfied please rate it thanks